OpenStudy (anonymous):

Andrew purchased a coat for $67.20 that regularly sells for $84.00. what was the percent discount that Andrew received?

OpenStudy (anonymous):

is it F.)16.8% G.)20% H.)25% I.)80%

OpenStudy (anonymous):

84-67.20=$16.8 difference 16.8/84 x 100 = 20% discount

OpenStudy (anonymous):

\[\frac{16.8}{84}\times100 = 20 %\]

OpenStudy (anonymous):

$16.8 is the difference calculated and we have to put it over the original amount of $84 to calculate the 20% discount received.

OpenStudy (anonymous):

how did you get 100? i feel that you got the right answer i just dont know where you got the 100 from?

OpenStudy (anonymous):

you see the question says what percent discount? percent you always have to times 100.

OpenStudy (anonymous):

16.8/84 = 0.2 and if they said in the question they wanted it in ratios than 0.2 is fine but they want it in percentage so 0.2 x 100= 20%

OpenStudy (anonymous):

oka.. thank you so much:)

OpenStudy (anonymous):

no problem, medal thanks :P

OpenStudy (anonymous):

your welcome:)

OpenStudy (anonymous):

I hope we can be friends and thanks once again.

OpenStudy (anonymous):

inbox me oka.?